• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ulmer

    Log entry time 16:02:08 on June20,2002

    Entry number 86506

    keyword=Shift summary

    Start shift on kinematics Q2D_g40 with 90+ uA beam.

    08:55 ROC 14 crash. Resetting crates.
    09:05 Back up and running.
    09:15 We tested (again) the Left bogie control and again the GUI would
    not pop up. So rebooted the crate and now the GUI pops up and we are
    able to enable the PS, etc. We are not ready to actually move the Left
    spectrometer for some time, but just wanted to make sure it would work
    when needed.
    10:10 ROC 14/1/2/3 problems. Resetting didn't work. Communication
    failures rampant. "Domain not bound" errors, ssh not working ...
    Bob Michaels is on his way to troubleshoot.
    10:20 MCC informs us that there are current trips and we will be getting pulsed
    10:30 MCC tells us we will be now getting some CW beam.
    10:55 Beam is back at 100 uA. We (Bob) are still dealing with DAQ problems
    11:10 Ask for controlled access to reset network switch. No effect.
    11:30 Out of controlled access.
    12:15 DAQ problem solved with help of Bob and Computer Center. Both adaqep
    and adaqcp, which are yp slave-servers, required rebooting. Problems with
    these machines gave problems with the Sun workstations which then
    made rebooting the ROC's impossible.
    12:20 Restart acquisition at 100 uA.
    15:45 MCC says they lost gun voltage and North linac dropped from beam
    permit to power permit with no indicated fault. Beam should be back
    15:50 Beam back.

    Run Summary for all good Q2D_g40 runs so far (goal = 400 counts)

    Run Events Counts (running total)

    1684 pointing run (previous shift)
    1685 0.57M events (previous shift) - 33 counts
    1686 0.48M events (previous shift) - 55 counts
    1687 1M events (in progress at shift start) - 110 counts
    1690 1M events - 150 counts
    1693 1M events - 195 counts
    1694 1M events - 234 counts
    1695 1M events - 279 counts
    1696 In progress at end of this shift.

    We get about 40 trues into all cuts per 1M events and 1M events takes about
    1 hour of steady 100 uA running. Given we want 400 counts total, this amounts
    to about 9-10 hours, instead of the 5.5 hours predicted. We started production
    running at about 06:30 (previous shift). The prediction of 5.5 hours would mean
    that with 60% efficiency we expected to finish this point at 15:30 or so.

    We could, instead, take 75% of the 400 counts = 300 counts and finish at about 16:30. Let's finish at 17:00 with this point (including efficiency measurement).
    At this time, Bodo can begin the ep and we will change setting to Q2D_g50.
    The arc energy measurement is deferred until tomorrow when we change settings
    again (either after Q2D_g50 or after Q2D_i20).